Solve for the unknown using the SCI format. Show your solution.
(Case A)
Expense - 20,000
Net Income/(Net Loss) - 50,000
Service Income - ?

(Case B)
Net Income/(Net Loss) - 67,500
Service Revenue - 80,000
Expenses - 12,500
Other Comprehensive Income - 10,500
Comprehensive Income - ?

(Case C)
Comprehensive Income - 40,000
Other Comprehensive Income - 4,000
Service Income - 130,000
Net Income/(Net Loss) - 36,000
Expenses - ?
1
Expert's answer
2021-09-07T08:31:12-0400

Solution:

Case A:

Service Income = Net Income (Net loss) + Expense

= 50,000 + 20,000

= 70,000


Case B:

Comprehensive Income = Net Income (Net Loss) + Other Comprehensive Income

= 67,500 + 10,500

= 78,000


Case C:

Expense = Service Revenue - Net Income (Net Loss)

= 130,000 - 36,000

= 94,000